OpenStudy (anonymous):

A dam needs a supporting beam. The dam leans at an 80 degree angle and is 200 feet tall. If the base of the supporting beam is placed 75 feet from the base of the dam and the beam extends to the top of the dam, how long must the beam be?

OpenStudy (anonymous):

Do you think you can draw this problem?

OpenStudy (anonymous):

and post it?

OpenStudy (anonymous):

is it this?

OpenStudy (anonymous):

Do you know the law of cosines? Well it says that:\[c^2=a^2+b^2-2abcos \theta\] c is the side opposite the angle:\[\theta\] and a and b are the sides adjacent to the angle In this case c is x, Ø is 80 degrees,a is 200, and b is 75 so plug that into the law of cosines to solve for x.

OpenStudy (anonymous):

so \[x^2=(200)^2+(75)^2-2•200•75•\cos(80º)\]

OpenStudy (anonymous):

so solve for x

OpenStudy (anonymous):

remember the law of cosines @Bella_Q